OpenStudy (anonymous):

Find the greatest common factor of -30x^4 yz^3 and 75x^4 z^2.

OpenStudy (whpalmer4):

\[GCF(-30x^4yz^3,75x^4z^2)\] Okay, you can do this piece by piece and multiply the components. What is the GCF of 30 and 75?

OpenStudy (anonymous):

15

OpenStudy (whpalmer4):

Good! What's the GCF of \(x^4\) and \(x^4\)? :-)

OpenStudy (anonymous):

x^4

OpenStudy (whpalmer4):

Correct again. Now the \(y\) portion only appears in 1 of the items, so it doesn't contribute to the GCF. What is the GCF of \(z^3\) and \(z^2\)?

OpenStudy (anonymous):

z^2

OpenStudy (whpalmer4):

Correctamundo. So now we multiply all of the components we found together. What's the answer?

OpenStudy (anonymous):

15x^4z^2

OpenStudy (whpalmer4):

That's correct. Another way you could think of this is factoring. Say we had \[-30x^4 yz^3 + 75x^4 z^2\]and wanted to write it in the form \(a(b+c)\). What could we factor out to use as \(a\)? That's the GCF! We could rewrite it as \[15x^4z^2(-2yz+5)\]right?

OpenStudy (anonymous):

Yes. I use factoring a lot in this unit of algebra:)

OpenStudy (whpalmer4):

If you want a more formal description, to find the GCF, we factor each of the items into their prime components: \[-30x^4yz^3= -1*2*3*5*x^4*y*z^3\]\[75x^4z^2 = 3*5^2*x^4*z^2\] Now we collect the highest common powers of each of the common factors and multiply them together: \[GCF = 3*5*x^4*z^2 = 15x^4z^2\]